7. Using the formula at the beginning of this lesson, calculate the wavelength of a signal from KPhET 98.7. Remember that FM stations transmit in the Megahertz range. This means that KPhET is transmitting a frequency of 9.87 x107 hz. What is the wavelength of the signal from this radio station?




8. My favorite radio station is NPR, which transmits a signal that is has a wavelength of 3.38 m. What is the frequency of this signal? Remember, light speed is 3 x108 m/s.




9. AM 630 is where I listen to sports. The frequency is 630 khz (or 630,000 hz). What is the wavelength?




10. Pick your favorite radio station, or one your friends listen to. What is its frequency and wavelength?


1
Expert's answer
2022-04-18T08:50:30-0400

7.

"\lambda=\frac c\nu=3.04~m ."

8.

"\nu=\frac c\lambda=8.88\cdot 10^7~Hz."

9.

"\lambda=\frac c\nu=476.2~m."